The time required to produce the first chair is 1701 minutes. Determine the time required to produce the 450th chair if the learning-curve coefficient is 0.85.
Answer #1

Y = a * X ^ b

Y = 1701 * 450 ^(log 0.85 / log 2)

Y = 1701 * 450 ^ (-0.234465)

Y = 1701 * 0.238733

Y = 406.09 minutes